rssLink RSS for all categories
 
icon_red
icon_green
icon_red
icon_red
icon_blue
icon_green
icon_green
icon_red
icon_red
icon_red
icon_orange
icon_green
icon_green
icon_green
icon_green
icon_blue
icon_green
icon_orange
icon_red
icon_green
icon_red
icon_red
icon_green
icon_red
icon_red
icon_red
icon_red
icon_orange
icon_green
 

FS#2747 — FS#6707 — Dallas,TX/Los Angeles,CA

Attached to Project— Network
Incident
Whole Network
CLOSED
100%
We had a breakdown on the lienk between Dallas,TX and Los Angeles,CA. This has been repaired within about 9 hours.

UTC: 2012-05-08 17:40
A fault in Dallas, USA is detected, we will start to investigate this issue and keep you updated with more information as it comes available.

UTC: 2012-05-08 18:21
Our fiber provider has confirmed that there is a cable cut between Dallas and Huston. No estimated time of repair given at this point.

UTC: 2012-05-08 19:26
Our fiber provider has informed that they have located the fault and that they have engaged a splicing team. There is no estimated time of arrival of the splicing team at this point. We will keep you updated of the progress until the issue is resolved.

UTC: 2012-05-08 20:49
The Outside Plant Manager at the location of the fiber cut advised that the splice crew has arrived. The physical location of the fiber break has been identified, and Field Services reported that it was caused by an auger. There is still no ETR given.

UTC: 2012-05-08 22:34
It has been informed that the fiber cut is located in Midlothian, TX, which is between Dallas, TX and Waco, TX, which is impacting unprotected services, including dark fiber.
The Outside Plant Manager reported that excavation is underway to break through nearly seven feet of rock where the new hand hole will be placed.
Protected services were restored after traffic was moved to a diverse path.
There is still no ETR given.

UTC: 2012-05-09 00:40
Please be informed that our Field Services advised that excavation efforts are still underway.
There is still no ETR given.

UTC: 2012-05-09 02:15

Our provider informed that splicing is complete and all damage is repaired at the North End of the fiber. However, the South facing cable was stuck in the pipe.
The field technician on site reported that the excavation crew had a difficulty to find the South End, but the excavation is now complete and they are in the process of pulling the cable through the hand hole at this time. It was informed that splicing will commence at approximately 02:00 GMT, but no ETTR has been provided.

UTC: 2012-05-09 03:43
Please be informed that our local provider has informed us that the traffic has been restored at 03:15 UTC. We will provide you with official Reason For Outage in 3 business days.

UTC: 2012-05-09 03:45
Our local provider informed that the traffic was restored at 03:15 UTC. We will provide you with official Reason For Outage report in 3 business days.

Date:  Thursday, 10 May 2012, 07:39AM
Reason for closing:  Done